# Directional Reasoning Trajectory Change (DRTC): Identifying Critical Trace Segments in Reasoning Models

### v1

@@ Abstract
Understanding how language models carry out long-horizon reasoning remains an open challenge. Existing interpretability methods often highlight tokens or spans correlated with an answer, but they rarely reveal where the model makes consequential reasoning turns, which earlier context causally triggers those turns, or whether the highlighted text actually steers the reasoning process. We introduce Directional Reasoning Trajectory Change (DRTC), a process-causal framework for interpreting long-form reasoning from a single on-policy rollout. DRTC first detects pivot decision points using uncertainty and distribution-shift signals, then applies receiver-side interventions that preserve the realized rollout without resampling the continuation while blocking information flow from selected earlier chunks only at a pivot. It quantifies whether each intervention redirects the direction of the model’s log-probability trajectory relative to the realized rollout direction, producing a signed per-chunk attribution score. We additionally compute turning-angle curvature changes on raw logits as a complementary diagnostic and introduce curvature signatures to summarize shared intervention-response geometry via curvature role paths. Empirically, DRTC acts like a “steering map” for long traces, yielding a compact ranked shortlist of influential chunks: across four reasoning models spanning multiple backbone families, directional influence is sharply concentrated (per-example $|\mathrm{DRTC}|$ shares yield Gini $\approx 0.50$$0.58$ and top-$5\%$ mass $\approx 0.23$$0.28$), and learned pivots induce stronger intervention magnitudes than matched random spans. In a scaling study on 500 MATH problems with R1-Distill-Qwen-1.5B, learned spans continue to outperform matched random spans (median $\Delta=0.409$, 355/500 positive; sign test $p=2.3\times 10^{-21}$). Overall, DRTC provides a causally grounded, trajectory-level view of how specific context elements steer reasoning under on-policy dynamics.
@@ 1 Introduction
To enable reviewer-auditable inspection, each run exports full provenance artifacts grounded in the model’s exact tokenized input. These include: (i) per-example summary tables with pivot statistics and attribution aggregates, (ii) per-configuration ranked chunk lists with stable chunk identifiers, (iii) integrity-checked mappings between chunk indices and token spans, and (iv) a self-contained interactive HTML reviewer interface that links raw text, pivot locations, masking interventions, and trajectory visualization (Figure LABEL:fig:drtc_ui_demo).
All reported DRTC and curvature quantities are reproducible directly from the exported JSON bundles without re-running model inference. This design allows reviewers to verify pivot discovery, intervention construction, per-pivot effects $(\delta_{k,i},w_{k,i})$, and final aggregation mechanics step-by-step on the exact realized rollout.
In practice, the signed per-chunk scores are sharply concentrated, yielding a compact, auditable shortlist of context segments to inspect within long traces and a concrete, mechanically grounded starting point for subsequent circuit-level or mechanistic analyses.
@@ 2 Related work
Recent interpretability work has shifted from single-pass explanations toward causal analysis of multi-step reasoning traces. Zhao et al. (2026) formalize decorative thinking, showing that only a small fraction of chain-of-thought steps exert meaningful causal influence; their True Thinking Score (TTS) estimates causal contribution via stochastic perturbation and resampling. Macar et al. (2025) similarly argue that a single realized trace is insufficient and use on-policy resampling to study the resilience of semantic content across trajectories. Complementary work models causal influence across autoregressive generations: Walker and Ewetz (2025) construct attribution graphs that quantify dependence on the prompt and prior outputs while preserving causal structure. In parallel, geometric approaches treat reasoning as a trajectory in representation space: Zhou et al. (2025) relate logical structure to trajectory geometry and analyze discrete turning behavior, while Curved Inference measures residual-stream bending under a pullback semantic metric derived from the unembedding map (Manson, 2025). Other studies analyze logit and layer-wise dynamics for interpretability and decoding, including DeLTa’s logit-trajectory prediction strategy for decoding (He et al., 2025) and stage-wise analyses of arithmetic signal progression across layers (Yan, 2025).
DRTC targets a distinct explanatory object: which earlier context chunks causally steer a realized long-horizon reasoning trace under on-policy generation. It uses pivot-local, deterministic receiver-side interventions and a directional trajectory target in log-probability space, with curvature used only as a complementary diagnostic of intervention-induced reorientation. A detailed comparison to resampling-based methods, attribution graphs, geometry-first analyses, and layer-wise dynamics appears in Appendix C (Table 17).
@@ 3.3 Pivot discovery and weighting
where $H_{t}$ is entropy, $\mathrm{margin}_{t}$ is the top-2 probability margin, and $S_{t}$ is a local distribution-shift signal measured by Jensen–Shannon divergence between token windows before and after $t$. We use $(\gamma_{H},\gamma_{M},\gamma_{S})=(1.0,1.0,0.5)$.
@@ 3.4 Receiver-side interventions and relevance screening
At each pivot $\tau_{k}$, we test whether an earlier chunk $c_{i}$ is causally relevant by applying a receiver-side intervention that blocks information flow from $c_{i}$ only at $\tau_{k}$. Concretely, we implement this as attention-edge masking at the prediction position, leaving the realized prefix unchanged while testing which earlier positions can influence the pivot distribution; similar edge-level and activation-patching interventions are standard tools in mechanistic interpretability (Geva et al., 2023; Conmy et al., 2023; Zhang and Nanda, 2024).
Let $\ell_{t}(\cdot)$ denote pre-softmax logits. We define a pivot-local screening effect on the baseline top token $y_{\tau_{k}}^{\star}$:
@@
where (i,k)(-i,k) denotes masking chunk $c_{i}$ at pivot $\tau_{k}$. Intuitively, $a_{k,i}$ measures whether removing $c_{i}$ weakens the pivot decision.
We map screening effects to a bounded relevance gate
@@
where $\hat{a}_{k,i}$ is a robustly scaled version of $a_{k,i}$ and $\beta_{k}$ is a pivot-specific calibration parameter (chosen from the median magnitude of nonzero effects). The gate downweights chunk–pivot pairs that do not reliably affect the pivot decision while preserving ordering among stronger effects.
@@ 3.5 Directional trajectory change and DRTC aggregation
Positive $\delta_{k,i}$ indicates that chunk $c_{i}$ supports the realized rollout direction at pivot $k$. For stability, we robustly normalize $\delta_{k,i}$ within each pivot (median/MAD scaling with clipping; details in Appendix).
@@
We treat $\mathrm{DRTC}(i)$ as the primary attribution score; auxiliary diagnostic decompositions are used only for internal checks and are reported in the Appendix.
3.6 Curvature diagnostics, signatures, and similarity
In addition to directional alignment, we compute a geometric diagnostic of how interventions change local turning behavior. Curvature is computed in raw logit space. For consecutive logits $(z_{t-1},z_{t},z_{t+1})$, define
@@ 3.6 Curvature diagnostics, signatures, and similarity
with standard clipping and small-$\varepsilon$ stabilization in implementation. For each pivot $\tau_{k}$ and chunk $c_{i}$, we measure the curvature change induced by the intervention:
@@
CurvImpact is diagnostic only: it describes intervention-response geometry and is not used as a causal attribution score.
We summarize curvature response patterns with a curvature signature vector
@@
and define curvature similarity between chunks $i$ and $j$ as $\langle s_{i},s_{j}\rangle$. These quantities are used to group chunks into structural roles.
3.7 Curvature role paths: structural reorientation summaries
To compactly summarize long rollouts, we introduce curvature role paths, a diagnostic synopsis of major structural reorientations. We construct a graph whose nodes are chunks and whose edges connect pairs with curvature similarity above a fixed threshold; connected components define roles. Roles are purely structural groupings defined by curvature signatures and do not depend on DRTC or screening.
For each pivot $\tau_{k}$, we score the curvature reorientation contributed by a role $r$ as
@@ 3.7 Curvature role paths: structural reorientation summaries
select the role with maximal $\mathrm{RoleImpact}(r,k)$, and choose a representative chunk with maximal $|\Delta\kappa_{k,i}|$ within that role. Ordering these dominant roles by pivot index yields a pivot-aligned curvature role path.
To distinguish structurally prominent but causally irrelevant reorientations from those that support the realized direction, we optionally annotate each role with aggregated DRTC:
$\mathrm{DRTC}(r)\;=\;\sum_{c_{i}\in r}\mathrm{DRTC}(i).$
This overlay is descriptive and does not affect role construction or pivot assignment. Overall, curvature role paths provide a compact qualitative summary of intervention-induced reorientation types that complements chunk-level DRTC attribution rather than replacing it.
@@ 4 Results
We evaluate DRTC on a fixed 24-example slice from the math-rollouts problems, using the dataset only for problem statements (Bogdan et al., 2025). For each example, the model generates an on-policy autoregressive reasoning rollout (greedy decoding with fixed seeds), segmented into fixed-stride chunks for attribution; prompt text and ground truth are retained for auditability but excluded from pivot discovery, interventions, scoring, and qualitative ranking. Unless otherwise stated, we use $K=8$ pivots per example.
We report results for four open-weight reasoning models: LFM2.5-1.2B-Thinking, Ministral-3B-Reasoning, Phi-4-Mini-Reasoning, and R1-Distill-Qwen-1.5B (AI, 2025; Liu et al., 2026; Peng et al., 2025; Guo et al., 2025), evaluated on the same 24 examples with identical decoding settings and seeds. We compare C0 (DRTC only), C8 (pivot curvature diagnostics), and C9 (matched random-span control with identical masking and scoring logic).
4.1 Attribution invariance under curvature diagnostics
@@
C0 and C8 differ only in whether curvature diagnostics are computed and logged; curvature is never used in pivot discovery, intervention construction, gating, or DRTC aggregation. We test invariance by computing, for each example, the Spearman rank correlation between per-chunk DRTC scores from C0 and C8. Across all four models (24 examples each), $\rho=1.000$ for every example (median $1.000$, 95% CI $[1.000,1.000]$), with identity scatter and zero score differences within floating-point tolerance. Figure 2 shows a representative identity scatter for R1-Distill-Qwen-1.5B; full per-model plots and auxiliary checks appear in Appendix A.1.
@@
Figure 2: Curvature invariance under diagnostic logging (representative model: R1-Distill-Qwen-1.5B). Per-chunk DRTC scores from C0 and C8 lie on the identity line ($\rho=1.000$), confirming that enabling curvature computation is strictly diagnostic and does not alter attribution.
4.2 DRTC attribution is concentrated and structured
@@
Across all models and configurations, DRTC attribution is concentrated rather than diffuse: a minority of chunks accounts for a disproportionate share of total directional influence. Using per-example normalized $|\mathrm{DRTC}|$ mass shares, median Gini coefficients range from 0.500.500.580.58, and the top $5\%$ of chunks account for roughly 0.230.230.280.28 of total influence. Cross-model histograms, concentration curves, and bootstrap confidence intervals are provided in Appendix A.2 (see Fig. 6).
4.3 Pivot weighting is non-redundant
We test whether pivot importance weighting ($u_{k}$) meaningfully affects attribution by recomputing DRTC scores post hoc with uniform pivot weights. Specifically, using the same per-pivot quantities $(\delta_{k,i},w_{k,i})$ obtained from C0/C8/C9, we recompute
$\mathrm{DRTC}_{\text{uniform}}(i)=\sum_{k}\tfrac{1}{K}\,w_{k,i}\,\delta_{k,i},$
replacing learned $u_{k}$ with $1/K$. Because $u_{k}$ enters only in the final aggregation and does not influence $\delta_{k,i}$ or $w_{k,i}$, this procedure is mathematically equivalent to rerunning the method with uniform pivot weights.
Across all models, Spearman correlations between weighted and uniform rankings remain high ($\rho\approx 0.97$$0.99$), but strictly less than 1. Equivalent ablations for all four models, including weighted-vs-uniform scatter plots and per-model summary statistics, are provided in Appendix A.6.
Correlations are high but strictly below 1.000, indicating that pivot importance weighting preserves the dominant attribution structure while introducing non-trivial reshuffling of chunk rankings. This contrasts with the C0 vs. C8 comparison (Section 4.1), where correlations are identically 1.000, confirming that curvature logging is purely diagnostic whereas pivot weighting modulates attribution aggregation.
4.4 Random-span falsification: intervention magnitude
@@
We evaluate whether learned pivots (C8) produce meaningfully stronger interventions than matched random spans (C9). As the primary falsification signal, we compute the per-example mean absolute pivot-local intervention magnitude, $\mathbb{E}_{k}[|\delta_{k,i}|]$. Across all four models, C8 exhibits strictly higher median per-example intervention magnitude than C9, with median gaps (C8 - C9) ranging from +0.039+0.039 to +0.178+0.178 and 95% bootstrap confidence intervals shown in Figure 3. C8 exceeds C9 for the majority of examples (see Appendix A.3). This effect is consistent across examples rather than driven by a small number of outliers. Additional rank-stability diagnostics (e.g., $\delta$-vs-full Spearman) appear in Appendix A.3.
@@
4.5 Curvature provides a complementary geometric diagnostic
Within C8, curvature is a complementary geometric diagnostic: across models, $|\mathrm{CurvImpact}|$ is positively associated with $|\mathrm{DRTC}|$, while signed $\mathrm{CurvImpact}$ has near-zero association with signed $\mathrm{DRTC}$. This is consistent with curvature tracking reorientation intensity rather than directional steering, and it is not used in pivots, gating, or attribution. Aggregate statistics and visual diagnostics are in Appendix A.4.
4.6 Adjacency and shuffle controls

### v2

@@ Abstract
Understanding how language models carry out long-horizon reasoning remains an open challenge. Existing interpretability methods often highlight tokens correlated with an answer, but rarely reveal where consequential reasoning turns occur, which earlier context triggers them under causal intervention, or whether highlighted text actually steers the rollout. We introduce Directional Reasoning Trajectory Change (DRTC), a process-causal method that (i) detects pivot decision points via uncertainty and distribution-shift signals and (ii) applies receiver-side interventions that preserve the realized continuation without resampling while blocking information flow from selected earlier chunks only at a pivot. DRTC measures how each intervention redirects the log-probability trajectory relative to the realized rollout direction, yielding signed per-chunk attributions; we also compute logit-space curvature changes and curvature signatures as a complementary geometric diagnostic. Across four reasoning models, influence is sharply concentrated (Gini $\approx 0.50$$0.58$, top-$5\%$ mass $\approx 0.23$$0.28$), and learned pivots induce stronger effects than matched random spans. In a 500-problem MATH scaling study with R1-Distill-Qwen-1.5B, learned spans continue to outperform matched random spans (median $\Delta=0.409$, 355/500 positive; $p=2.3\times 10^{-21}$), and curvature-impact co-localizes with DRTC within traces (diagnostic). We benchmark against gradient- and perturbation-based chunk attributions and show graded outcome linkage: under embedding-interpolation edits, top-ranked DRTC chunks reduce teacher-forced gold-answer log-probability more than strict position-matched random chunks on a stability-filtered subset. Overall, DRTC provides a causally grounded view of how specific context elements steer on-policy reasoning trajectories.
@@ 1 Introduction
To enable reviewer-auditable inspection, each run exports full provenance artifacts grounded in the model’s exact tokenized input. These include: (i) per-example summary tables with pivot statistics and attribution aggregates, (ii) per-configuration ranked chunk lists with stable chunk identifiers, (iii) integrity-checked mappings between chunk indices and token spans, and (iv) a self-contained interactive HTML reviewer interface that links raw text, pivot locations, masking interventions, and trajectory visualization (Figure LABEL:fig:drtc_ui_demo). All reported DRTC and curvature quantities are reproducible directly from the exported JSON bundles without re-running model inference. This design allows reviewers to verify pivot discovery, intervention construction, per-pivot effects $(\delta_{k,i},w_{k,i})$, and final aggregation mechanics step-by-step on the exact realized rollout. In practice, the signed per-chunk scores are sharply concentrated, yielding a compact, auditable shortlist of context segments to inspect within long traces and a concrete, mechanically grounded starting point for subsequent circuit-level or mechanistic analyses.
@@ 2 Related work
Recent interpretability work has shifted from single-pass explanations toward causal analysis of multi-step reasoning traces. Zhao et al. (2026) formalize decorative thinking and propose True Thinking Score (TTS), estimating step-level causal contribution via stochastic perturbation and resampling; Macar et al. (2025) similarly use on-policy resampling to study semantic resilience across trajectories. Complementary lines of work model influence across autoregressive generations with attribution graphs (Walker and Ewetz, 2025), analyze reasoning as a geometric trajectory in representation space (Zhou et al., 2025; Manson, 2025), and study logit/layer-wise dynamics for interpretability and decoding (He et al., 2025; Yan, 2025).
[…]
@@ 3.3 Pivot discovery and weighting
where $H_{t}$ is entropy, $\mathrm{margin}_{t}$ is the top-2 probability margin, and $S_{t}$ is a local distribution-shift signal measured by Jensen–Shannon divergence between token windows before and after $t$. We use $(\gamma_{H},\gamma_{M},\gamma_{S})=(1.0,1.0,0.5)$. We compute $S_{t}$ as the Jensen–Shannon divergence between the mean next-token distributions (averaged softmax probabilities over $\mathcal{V}$) in a $W$-token window before $t$ (indices [tW,t)[t\!-\!W,t)) and a $W$-token window after $t$ (indices (t,t+W](t,t\!+\!W]), excluding position $t$ from both windows (default $W=5$).
@@
Defaults. Unless otherwise stated, we use $K=8$ pivots, enforce a minimum pivot spacing of 4 tokens to avoid near-duplicate pivots, and compute the JSD shift signal using 5-token windows before and after $t$; sensitivity to these choices is reported in Appendix A.7.
@@ 3.4 Receiver-side interventions and relevance screening
[…]
Attention-edge masking at a pivot.
Let chunk $c_{i}$ correspond to token indices $\mathcal{I}_{i}\subset\{1,\dots,\tau_{k}-1\}$ in the realized prefix. For each transformer layer $\ell$ and query head $h$, let $s^{(\ell,h)}_{\tau_{k},j}$ denote the pre-softmax attention score from query position $\tau_{k}$ to key position $j$. Our receiver-side mask sets
$s^{(\ell,h)}_{\tau_{k},j}\leftarrow-\infty\quad\text{for all }j\in\mathcal{I}_{i},$ (4)
[…]
Pivot-local screening effect.
Let $\ell_{t}(\cdot)$ denote pre-softmax logits at position $t$. We define a pivot-local screening effect on the baseline top token $y_{\tau_{k}}^{\star}$:
@@
where (i,k)(-i,k) denotes masking chunk $c_{i}$ only at pivot $\tau_{k}$ as above. Intuitively, $a_{k,i}$ measures whether removing access to $c_{i}$ weakens the pivot decision in the realized rollout context.
Relevance gate.
We map screening effects to a bounded relevance weight
@@
where $\hat{a}_{k,i}$ is a robustly scaled version of $a_{k,i}$ (e.g., median/MAD robust $z$-scoring with clipping) and $\beta_{k}$ is a pivot-specific calibration parameter chosen from a robust statistic of nonzero screening magnitudes at pivot $\tau_{k}$. The gate downweights chunk–pivot pairs that do not reliably affect the pivot decision while preserving ordering among stronger effects.
@@ 3.5 Directional trajectory change and DRTC aggregation
Pivot-local readout. Although $g$ is defined globally from the pivot endpoints, the intervention effect is read out locally at each pivot via $e_{k,i}=s_{\tau_{k}}-s_{\tau_{k}}^{(-i,k)}$ and $\delta_{k,i}$; DRTC does not rely on a final-pivot-only score. Alternative definitions of $g$ and meander robustness are reported in Appendix A.2. Positive $\delta_{k,i}$ indicates that chunk $c_{i}$ supports the realized rollout direction at pivot $k$. For stability, we robustly normalize $\delta_{k,i}$ within each pivot (median/MAD scaling with clipping; sensitivity and implementation details are in Appendix A.7).
@@
We treat $\mathrm{DRTC}(i)$ as the primary attribution score; auxiliary diagnostic decompositions and sanity checks, including pivot-weight ablations and per-component analyses, are reported in the Appendix (Appendix A.12 and Appendix A.7).
3.6 Curvature diagnostics and role paths
In addition to directional alignment, we compute a geometric diagnostic of how interventions change local turning behavior in raw logit space. For consecutive logits $(z_{t-1},z_{t},z_{t+1})$, define
@@ 3.6 Curvature diagnostics and role paths
with standard clipping and small-$\varepsilon$ stabilization in implementation. For each pivot $\tau_{k}$ and chunk $c_{i}$, we measure the intervention-induced curvature change
@@
$\mathrm{CurvImpact}$ is diagnostic only: it summarizes intervention-response geometry and is not used as a causal attribution score.
Curvature signatures and similarity.
We summarize curvature response patterns with a normalized signature
@@
and define curvature similarity between chunks $i$ and $j$ as $\langle s_{i},s_{j}\rangle$.
Roles and curvature role paths.
To compactly summarize long rollouts, we form an undirected graph over chunks where an edge connects $(i,j)$ if $\langle s_{i},s_{j}\rangle$ exceeds a fixed threshold; connected components define curvature roles $r$. Roles are purely structural groupings defined by curvature signatures and do not depend on DRTC or screening. For each pivot $\tau_{k}$, we score the reorientation attributed to role $r$ by
@@
select the maximizing role $r_{k}^{\star}=\arg\max_{r}\mathrm{RoleImpact}(r,k)$, and choose a representative chunk $i_{k}^{\star}=\arg\max_{c_{i}\in r_{k}^{\star}}|\Delta\kappa_{k,i}|$. Ordering the dominant roles $(r_{k}^{\star})_{k=1}^{K}$ by pivot index yields a pivot-aligned curvature role path. Optionally, we annotate each role with aggregated directional attribution
$\mathrm{DRTC}(r)\;=\;\sum_{c_{i}\in r}\mathrm{DRTC}(i),$
as a descriptive overlay (it does not affect role construction or pivot assignment). Overall, curvature role paths provide a compact synopsis of intervention-induced reorientation types that complements chunk-level DRTC attribution.
@@ 4 Results
We evaluate DRTC on a fixed 24-example slice from math-rollouts (Bogdan et al., 2025). For each example, the model generates a greedy on-policy reasoning rollout (fixed seeds), segmented into fixed-stride chunks for attribution; prompt text and ground truth are retained for auditability but excluded from pivot discovery, interventions, and scoring. Unless otherwise stated, we use $K=8$ pivots per example. We report results for four open-weight reasoning models—LFM2.5-1.2B-Thinking, Ministral-3B-Reasoning, Phi-4-Mini-Reasoning, and R1-Distill-Qwen-1.5B (AI, 2025; Liu et al., 2026; Peng et al., 2025; Guo et al., 2025)—evaluated on the same examples with identical decoding settings. We compare C0 (DRTC only), C8 (DRTC + curvature diagnostics), and C9 (matched random-span control with identical masking and scoring). On R1-Distill-Qwen-1.5B with $K=8$ on the 24-example slice, median wall time is 12.28s per example with 4864MB peak GPU memory (Appendix A.5).
4.1 Core findings: invariance, concentration, and falsification
Curvature logging is strictly diagnostic.
@@ 4.1 Core findings: invariance, concentration, and falsification
C0 and C8 differ only in whether curvature diagnostics are computed and logged; curvature is never used in pivot discovery, intervention construction, gating, or DRTC aggregation. Across all four models (24 examples each), per-example Spearman correlations between C0 and C8 chunk rankings are identically $\rho=1.000$ (median $1.000$, 95% CI $[1.000,1.000]$). Figure 2 shows a representative identity scatter for R1-Distill-Qwen-1.5B; full invariance diagnostics appear in Appendix A.1.
@@
Figure 2: Curvature invariance under diagnostic logging (representative model: R1-Distill-Qwen-1.5B). Per-chunk DRTC scores from C0 and C8 lie on the identity line ($\rho=1.000$), confirming curvature computation is purely diagnostic.
DRTC yields a compact shortlist of influential chunks.
@@
Across models, directional influence is concentrated rather than diffuse: using per-example normalized $|\mathrm{DRTC}|$ mass shares, median Gini coefficients range from 0.500.500.580.58 and the top $5\%$ of chunks account for roughly 0.230.230.280.28 of total influence. Cross-model histograms, concentration curves, and bootstrap confidence intervals are in Appendix A.3.
Random-span falsification: learned pivots induce stronger interventions.
@@
As a falsification signal, we compare learned pivots (C8) to matched random spans (C9) under identical masking and scoring. Across all four models, C8 exhibits strictly higher median per-example mean pivot-local intervention magnitude $\mathbb{E}_{k}[|\delta_{k,i}|]$ than C9, with median gaps (C8-C9) ranging from +0.039+0.039 to +0.178+0.178 (95% bootstrap CIs in Figure 3); per-example diagnostics are in Appendix A.4.
@@
Pivot weighting is non-redundant.
Recomputing DRTC post hoc with uniform pivot weights ($u_{k}=1/K$) preserves the dominant structure but induces non-trivial rank reshuffling: weighted-vs-uniform Spearman correlations remain high ($\rho\approx 0.97$$0.99$) but strictly below $1.000$ across all models. Full ablations and scatter plots are in Appendix A.12.
4.2 Robustness, baselines, curvature calibration, and outcome linkage
We summarize additional quantitative evidence here and provide full tables/figures in the Appendix.
Robustness to direction estimation and heuristic knobs.
Chunk rankings are stable to alternate direction-estimation strategies for $g$ (Appendix A.2). A targeted sensitivity sweep shows near-identity agreement under $\beta$-gate rescaling (median Spearman vs. baseline =0.9901=0.9901 at $\beta$-scale $0.5$; $0.9892$ at $\beta$-scale $2.0$), while varying the pivot budget is more consequential than $\beta$ rescaling (median Spearman vs. baseline =0.3376=0.3376 at $K=4$ and $0.7298$ at $K=12$; baseline $K=8$). Additional knob sweeps and concentration metrics are reported in Appendix A.7.
Intervention specification and masking ablations.
Receiver-side attention-edge masking is applied at the pivot across all layers and query heads; ablations over head subsets (and related implementation details, including GQA/MQA handling) are reported in Appendix A.6.
Comparisons to attribution baselines.